You are given the following information The firm has 200,000 shares outstanding for the entire year. It had net income of 250,000. It also had 2,000 7%, 1,000 Par- value bonds convertible into 25 shares of common each. The firm's tax rate is 40%. The firm has no preferred stock. What is fully diluted EPS
Answer #1

Α Hmm Net income Add: Interest net of tax Adjusted Income 250000 84000 334000 6 7 9 10 11 No of shares Add: Potential equity

ovou wn Net income Add: Interest net of tax Adjusted Income 250000 =2000*1000*7%*(1-40%) =SUM(C4:05) No of shares Add: Potent